OpenStudy (angelwings996):

Describe the vertical asymptotes and holes for the graph of y=(x-5)/(x^2+4x+3)

OpenStudy (angelwings996):

\[y=\frac{ x-5 }{x ^{2} +4x+3}\]

OpenStudy (anonymous):

what do u mean by describe

OpenStudy (anonymous):

Factor the denominator: x^2 + 4x + 3 = x^2 + x + 3x + 3 = x(x + 1) + 3(x + 1) = (x + 3)(x + 1) Since the numerator and denominator have no common factors, there are no holes. To find vertical asymptotes, the denominator must be zero (and the numerator must be non-zero): (x + 3)(x + 1) = 0 x + 3 = 0 or x + 1 = 0 x = -3 or x = -1 The answer is asymptotes: x = –3, –1 and no holes.

OpenStudy (angelwings996):

Idk, that's what the problem asks but its asking what the answers are

OpenStudy (angelwings996):

Thank you! @seiga That's what I thought for the asymptotes but I wasn't sure on how to get the holes

OpenStudy (anonymous):

no problem(: